OpenStudy (anonymous):

a rectangle is 5 units more than the side of a square, and width half the side of the square. If the two areas are equal, what are the square's dimensions?

OpenStudy (saifoo.khan):

As their areas are same, we will write them as: A = b*h A = A b*h = b* h (x+5) * (0.5x) = x^2 0.5x^2+ 2.5x = x^2 -0.5x^2 + 2.5x = 0 therefore - factor them.. x = 0 & x = 5

OpenStudy (saifoo.khan):

So the dimentions are 5.

OpenStudy (anonymous):

thanks for showing your work, helps a lot!!

OpenStudy (saifoo.khan):

you welcome! :)